A Complete Bibliography of Proceedings of the ACM on Computer Graphics and Interactive Techniques

(PACMCGIT)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

25 May 2022 Version 1.02

Title word cross-reference

3 [DAI^{+}18, FLS^{+}22, FXX^{+}22, NF21, STSK20].

/Good [Rüs21].

60Hz [RW21].

Accelerate [WMZ^{+}20]. Accelerated [LVY^{+}20, KKSM18]. Accuracy [Wan18]. Accurate [KK19]. acoustic [CASR21]. Adaptive [EK18, MYS20, RHH^{+}22, SPD18, WSG19, XO21, YZK^{+}19, XOKN20]. Address [SLY20]. Advection [NZT19].

Advection-Reflection [NZT19]. Agent [HG19, YW19]. Agent-based [YW19]. agents [CASR21]. AI [CASR21]. Aiming [KKS^{+}20]. algorithm [WH21]. Algorithms [DXS21, HRZ^{+}19]. Aliasing [CLS^{+}21]. Alpha [Yuk18]. Altered [WH21]. Alternating [ANAM^{+}20]. Am [RW21]. Ambient [VSE21, SBE22]. Andean [MD21]. Animals [YBB^{+}21]. Animation [FLS^{+}22, YY21, YW19]. Anisotropic [KHDN22]. Anti [CLS^{+}21]. Anti-Aliasing [CLS^{+}21]. Apparent [LD18, RPP21]. Appearance [ANEK21]. application [Dup20]. Applications [CKY^{+}22]. Approach [RPP21, STSK20, XK21]. Approaches [FJdNJ22]. Approximate [RPHD20]. Approximation [LPW20, TYS20]. Approximations [STK20]. Architecture [JWB^{+}21]. Art [Fuj21, ZRL21]. Artistic [NF21]. Assistive [KLXvdP21]. Attribute [PKM22]. Audio [FLS^{+}22]. Audio-Text [FLS^{+}22]. Avatars [CO18]. Averaging [DXS21].

Bad [Rüs21]. Based [JDZD19, JFS^{+}21, LBR^{+}18, LBG18, MCC22, MKKP18, PJL21, RPHD20, RPP21, SHS^{+}21, SDY^{+}18, TKC21, XK21, ARM^{+}19, BWL18, FJdNJ22, KK19, Muk18, PLRD21, TTK^{+}21, WH21, YW19, YY21]. Better [RW21]. Binary [Dup20]. Biosignals [YBB^{+}21]. bisection [Dup20]. Blend [PKM22]. Blending [HN18]. Bodies [MWW18, NF21]. Body [LAJB18]. Bones [Muk18]. Botanical [QLZF21]. box [YW19]. box-manipulation [YW19]. Bringing [KHDN22]. Buffers [HCNS18]. Builder [VKJ^{+}18]. Building [KK19]. BVH [VKJ^{+}18]. By-Example [HN18].

Cache [KKI^{+}18]. Calculate [Wan18]. Can [Far21]. Cangjie [ZRL21]. Caps [PD19]. Capture [SPL^{+}21]. Cars [RPHD20]. centric [HRZ^{+}19]. Character [XK21, YY21]. Cheat [JFS^{+}21]. codec [KKSM18]. Coding [PKM22]. Collision [MEM^{+}20, TLTM18]. Communication [HG19]. Compacted [SLY20]. Companion [YBB^{+}21]. Comparative [SLW^{+}21]. Compensated [SGH18]. Competitive [JFS^{+}21]. Complicated [JWB^{+}21]. Compression [LMSS18, PKM22, SLY20, vdLSE20]. Computer [Sim21]. Concurrent [Dup20]. Conditions [KKS^{+}20]. Conductors [MD21]. Conical [XJZ^{+}22]. consistent [SBE22]. Constrained [RHH^{+}22]. Contact [HGG^{+}19, LB19]. Contacts [BBGB20]. Content [YZK^{+}19]. Continuity [MD21]. Control [BWL18, KBFF^{+}21, XO21, XK21].

Conversations [JDZD19]. Conversion [Muk18]. Convolution [JWB^{+}21, TKC21]. Convolutional [CAS22]. Cooperative [SSH^{+}21, YW19]. Coordinating [HG19]. Corotated [KBFF^{+}21]. Cosines [KHDN22]. Cost [MUEM22]. Coupling [ANEK21]. CPU [SLY20]. CPU/GPU [SLY20]. Create [FXX^{+}22]. Crowd [DMH^{+}21]. CSG [ZCL18]. Cubic [TYS20]. Culling [TLTM18, XJZ^{+}22]. Curves [TYS20].

D

[DAI^{+}18, FLS^{+}22, FXX^{+}22, NF21, STSK20]. Data [Day21, HMES20, KK19, SLY20, TKC21]. Data-Driven [TKC21]. DCGrid [RHH^{+}22]. Deep [CWZ^{+}18, DAI^{+}18, FJdNJ22, JDZD19, LWM19, SDY^{+}18, TKC21]. Deferred [MYS20]. Deformable [CDGB19, LWM19]. Deformer [BBGB20]. Denoising [HMES20]. Dependent [LBG18]. Depth [FHSS18, HCNS18, VSE21, KKSM18]. Design [BWL18, Far21, MWW18]. Designed [FJdNJ22]. Destruction [WFM21]. Detection [JFS^{+}21]. Difference [ANAM^{+}20]. Diffuse [dDB22]. Digital [HCR^{+}21, NF21]. Dimensional [CDGB19, QLZF21]. Disconnect [Day21]. Disk [WS18]. Displacement [LMSS18]. Displays [LD18]. Distance [MEM^{+}20]. Distribution [Yuk18]. Divergence [Wan18]. Diverse [HGG^{+}19, PJL21]. Domains [PJL21]. Driven [FLS^{+}22, TKC21]. drop [Fuj21]. Dual [LVY^{+}20]. Dual-Split [LVY^{+}20]. Ductile [WDG^{+}19]. Dynamic [JWB^{+}21]. Dynamics [CO18].

Easily [FXX^{+}22]. EasyVRModeling [FXX^{+}22]. edge [Dup20]. Effect [WSX^{+}22]. Effects [RW21, WSG19]. Efficient [AHGB19, CASR21, DXS21, MYS20, SCNW19, SGH18, YY21]. ejections [LRC^{+}22]. Elastic [ARM^{+}19, BBGB20, CDGB19, KBFF^{+}21, MWW18]. Elastic-rigid [BBGB20]. Electronic [MD21]. Elucidated [Sim21]. Embodied [BDWG21]. Energized [LAJB18]. Energy [KBFF^{+}21, SGH18]. Energy-Compensated [SGH18]. Enhancement [LD18]. Environment [WH21]. Environmental [XJZ^{+}22]. Error [CKY^{+}22]. Estimation [SPD18]. Evaluating [Wan18]. Evaluation [DMH^{+}21]. Evaluator [ANAM^{+}20]. Events [FJdNJ22]. Example [HN18, TTK^{+}21]. Example-based [TTK^{+}21]. Experience [HCR^{+}21, WSX^{+}22, WSG19, ZRL21]. Expressive [FLS^{+}22]. Extracted [Wan18]. Eyck [Sim21]. Eye [BDWG21, JDZD19]. Eye-Tracking [BDWG21].

Face [KK19]. FaceBlit [TTK^{+}21]. Faces [MUEM22]. Facial [FLS^{+}22, TTK^{+}21]. Fast [KBFF^{+}21, VKJ^{+}18, WS18, Wan18]. faster [GBW20]. Feature [MSK22]. Feature-Wise [MSK22]. Feedback [MCC22]. Feminism [Far21]. Field [FHSS18, MEM^{+}20, KKSM18]. Fields [CWZ^{+}18]. Figures [NF21]. Filter [TY18]. Filtering [CLS^{+}21, CAS22, SPD18]. Filters [STK20]. Flexible [KLXvdP21]. FLIP [ANAM^{+}20]. Flow [RPHD20, RPP21, WXCT19]. Flows [WFM21]. Fluid [RHH^{+}22, TY18, WXCT19, YCL^{+}19]. fly [KKT^{+}18]. Footage [FJdNJ22]. Forces [KLXvdP21]. Foveated [MDZV18]. Fracture [LAJB18, WDG^{+}19]. Framework [HRZ^{+}19]. Free [PLRD21]. Free-view [PLRD21]. Friction [ANEK21]. Frictional [HGG^{+}19]. functional [MMMC21]. Functions [Kir18].

G [RW21]. G-SYNC [RW21]. Game [FJdNJ22, SLW^{+}21, SSH^{+}21]. Gameplay [FJdNJ22, RW21, WSX^{+}22, WSG19]. Games [WSX^{+}22, YZK^{+}19]. Gaming [JFS^{+}21]. GAN [WXCT19, XK21]. GAN-Like [XK21]. General [HRZ^{+}19]. Generalized [Oga20]. Generation [JDZD19, LBG18, SDY^{+}18, WS18]. Generative [PJL21]. Genetic [WH21]. Genetic-algorithm-based [WH21]. Geometric [AHGB19, CLS^{+}21, WJG^{+}21]. Geometry [KKT^{+}18, QLZF21, vdLSE20]. GGX [KHDN22]. Glint [CLS^{+}21]. Global [SPL^{+}21]. Good [Rüs21]. GPU [KKI^{+}18, KKSM18, RHH^{+}22, SLY20]. GPU-accelerated [KKSM18]. GPUs [TLTM18]. Gradient [SPD18]. Graph [PJL21]. Graph-Based [PJL21]. Grid [LY19, RHH^{+}22]. Guided [KW21, XOKN20].

Hair [CAS22, RL18]. Haptic [MCC22].

Hardware

[LVY^{+}20, MYS20, VKJ^{+}18, WMZ^{+}20]. Hardware-Accelerated [LVY^{+}20]. Hashing [TLTM18, ZCL18]. HCSG [ZCL18]. Head [JDZD19, LD18]. Head-mounted [LD18]. Helper [Muk18]. Heterogeneous [PHM^{+}21]. HeterSkinNet [PHM^{+}21]. Hierarchical [HMN18]. Hierarchy [LY19]. High [HN18, KKS^{+}20, LMSS18, STK20, VKJ^{+}18, vdLSE20, KKSM18]. High-Performance [HN18, STK20]. High-Quality [VKJ^{+}18, KKSM18]. High-Resolution [LMSS18]. Histogram [HN18]. Histogram-Preserving [HN18]. Human [MUEM22, MSK22, NF21, ZRL21]. Human-Machine [ZRL21]. Hybrid [HGG^{+}19, PLRD21]. Hyperparameter [YY21].

Identify [FJdNJ22]. Image

[LBR^{+}18, PLRD21, RPHD20, STK20].

Image-Based

[LBR^{+}18, RPHD20, PLRD21]. Images [ANAM^{+}20]. Imaginary [LZY21]. Imitation [XK21]. Immersive [FXX^{+}22, HCR^{+}21]. Implicit [KBFF^{+}21]. Importance [EK18, XOKN20]. Improves [KKS^{+}20]. In-Depth [HCNS18]. Increase [RPP21]. Independent [MKKP18, BBGB20]. Indigenous [MD21]. Indoor [JWB^{+}21]. Information [WJG^{+}21]. Input [KKS^{+}20]. Insects [WH21]. Installation [Fuj21]. Instant [TTK^{+}21]. Instruments [BDWG21]. Interaction [SHS^{+}21]. Interactive [CO18, Fuj21, HHCM21, LRC^{+}22, MCC22, MWW18, SPL^{+}21, SDY^{+}18, XK21, ZRL21]. Interface [BDWG21]. Interfaces [dDB22]. Intersections [WMZ^{+}20]. Intersector [RL18]. Invariant [ARM^{+}19]. Investigating [FJdNJ22]. Invisible [Rüs21]. Isosurface [Wan18].

Jan [Sim21]. Joint [FLS^{+}22]. Julian [Day21].

Kernel [MDZV18].

Labels [RPHD20]. Labor [Rüs21]. Late [KKS^{+}20]. Latency [KKS^{+}20]. Layer [FHSS18]. Layered [dDB22]. Layouts [JWB^{+}21]. Learned [WJG^{+}21]. Learning [CO18, HG19, JDZD19, RPP21, TKC21, XK21, YW19]. Learning-Based [JDZD19]. Level [STSK20]. Light [CWZ^{+}18, Oga20, KKSM18]. Lightcuts [LY20]. Lighting [LY19, XJZ^{+}22]. Lights [EK18, LPW20]. Like [XK21]. Linearly [KHDN22]. Liquid [TKC21]. Liquids [AHGB19, RPP21]. Living [WH21]. Local [MEM^{+}20, MSK22]. Locomotion [MSK22]. LOD [LBG18]. Long [WMZ^{+}20]. longest [Dup20]. Lossless [YZK^{+}19]. Lossy [vdLSE20]. Low [MUEM22]. Low-Cost [MUEM22].

Machine [ZRL21]. Making [NF21].

Manifold [CDGB19]. manipulation [YW19]. Many [EK18]. Map [CLS^{+}21, SDY^{+}18]. Mapping [Fuj21, STSK20]. Mask [Far21]. masks [GBW20]. Massively [KKT^{+}18]. Massively-Parallel [KKT^{+}18]. Material [HGG^{+}19, WDG^{+}19]. Materials [HGG^{+}19, dDB22]. Measurement [LZY21]. Medical [HMES20]. Meditation [HCR^{+}21]. Memories [YBB^{+}21]. Memory [RHH^{+}22]. Memory-Constrained [RHH^{+}22]. Mesh [WS18]. Method [HGG^{+}19, Wan18, WDG^{+}19]. Methods [SLW^{+}21]. Metric [DMH^{+}21, LBG18, Wan18]. Mimicry [WH21]. Ming [HCR^{+}21]. Mitigation [SLW^{+}21]. Mixed [CDGB19, CNI^{+}20, SSH^{+}21]. Mixed-Dimensional [CDGB19]. Mixed-Reality [SSH^{+}21]. Mobile [CNI^{+}20]. Mode [KBFF^{+}21]. Model [CDGB19, FLS^{+}22, JDZD19, PJL21]. Modeling [LB19, NF21, ZCL18]. Modelling [MSK22]. Models [FXX^{+}22, KK19]. modern [KKI^{+}18]. Modified [SLY20]. Modulated [KLXvdP21]. Moment [MKKP18]. Moment-Based [MKKP18]. Mother [Rüs21]. Motion [JDZD19, KLXvdP21, MSK22, PJL21, SPL^{+}21, SCNW19, YZK^{+}19]. mounted [LD18]. Movement [LZY21]. Multi [DAI^{+}18, FHSS18, HG19, LB19, WXCT19]. Multi-Agent [HG19]. Multi-Layer [FHSS18]. Multi-Pass [WXCT19]. Multi-Resolution [LB19]. Multi-View [DAI^{+}18]. Multigrid [AHGB19]. Multiple [PJL21].

Narrow [TY18]. Narrow-Range [TY18]. Nature [WH21]. Navigation [HG19]. Network [PHM^{+}21]. Networks [CAS22, FJdNJ22, JWB^{+}21, SCNW19, SDY^{+}18]. Networks-based [FJdNJ22]. Neural [CAS22, FJdNJ22, HMES20, RPP21, SCNW19, SDY^{+}18]. Noise [Day21, HN18, Kir18]. Non [CDGB19, Kir18, Muk18]. Non-Manifold [CDGB19]. Non-periodic [Kir18]. Non-rigid [Muk18]. Nonlinear [CO18]. Normal [CLS^{+}21, SDY^{+}18]. Novel [HRZ^{+}19].

Objects [CDGB19, LWM19, GBW20]. Occlusion [VSE21, SBE22]. Oliver [Day21]. On-the-fly [KKT^{+}18]. opacity [GBW20]. Operator [HN18]. Optimization [KLXvdP21, MEM^{+}20, YY21]. Optimized [LBR^{+}18]. Optimizing [KKI^{+}18]. Order [MKKP18, NZT19]. Order-Independent [MKKP18].

Paglen [Day21]. Parallel [HRZ^{+}19, KKT^{+}18, TLTM18]. Particle [RPP21, TKC21]. Particle-Based [RPP21, TKC21]. party [JDZD19]. Pass [WXCT19, XOKN20]. Passthrough [CNI^{+}20]. Patch [LBG18]. Patch-Based [LBG18]. Path [HMES20, HHCM21]. Pattern [BWL18]. perception [CASR21]. Perceptually [DMH^{+}21]. Perceptually-Validated [DMH^{+}21]. Performance [FJdNJ22, HN18, STK20, WSX^{+}22, WSG19]. periodic [Kir18]. Permutation [PKM22]. Perspectival [Sim21]. Perturbations [WJG^{+}21]. Phantom [RL18]. Phases [MSK22]. Photometric [LPW20]. Photon [SGH18]. Physics [KK19, MCC22, XK21, YY21]. Physics-Based [MCC22, XK21, KK19, YY21]. Pixel [TKC21]. Player [SSH^{+}21]. Playing [RW21]. PLOCTree [VKJ^{+}18]. plume [LRC^{+}22]. Poetics [Rüs21]. Poetry [ZRL21]. Point [HGG^{+}19, WDG^{+}19]. Poisson [WS18]. Poisson-Disk [WS18]. Politics [Day21, Rüs21]. Polygonal [LPW20]. Portals [Oga20]. Position [ARM^{+}19, SPL^{+}21]. Position-based [ARM^{+}19]. Post [KKS^{+}20]. Post-Render [KKS^{+}20]. Precomputed [LWM19]. Predicting [CKY^{+}22]. Prediction [DAI^{+}18, PHM^{+}21, SPL^{+}21]. Preserving [HN18, SLP21]. Primitive [WMZ^{+}20]. Problems [DXS21]. Procedural [Kir18]. Processing [KKT^{+}18, KKI^{+}18]. Progressive [WS18]. Projected [PD19]. Projection [Fuj21, WS18]. Projection-Mapping [Fuj21]. Projective [KB18]. Proposition [Day21]. PSCC [TLTM18]. pyroclastic [LRC^{+}22].

Quadratic [TYS20]. Quality [DMH^{+}21, VKJ^{+}18, KKSM18].

Racing [SLW^{+}21]. Radiance [LWM19]. Range [TY18]. Ray [LMSS18, RL18, WMZ^{+}20, XJZ^{+}22, GBW20]. Ray-Hair [RL18]. Ray-Traced [XJZ^{+}22]. Ray/Primitive [WMZ^{+}20]. Real [CKY^{+}22, CNI^{+}20, CLS^{+}21, CAS22, LD18, LY19, LY20, LPW20, MUEM22, MSK22, PD19, SPD18, SCNW19, TTK^{+}21, WH21, XOKN20, XO21, XJZ^{+}22, YCL^{+}19, KKSM18, ZCL18]. Real-Time [CKY^{+}22, CLS^{+}21, CAS22, LY19, LY20, MUEM22, MSK22, XJZ^{+}22, YCL^{+}19, CNI^{+}20, LD18, LPW20, SPD18, SCNW19, TTK^{+}21, WH21, XOKN20, XO21, KKSM18, ZCL18]. Reality [CNI^{+}20, HMN18, SLW^{+}21, SSH^{+}21, WSX^{+}22, ZRL21]. Recall [YBB^{+}21]. Reconstruction [HHCM21, QLZF21, KKSM18]. Recovering [WJG^{+}21]. Redirected [SHS^{+}21]. Reflection [NZT19, RPHD20]. Region [LBR^{+}18]. reinforcement [YW19]. Relighting [MUEM22]. ReMember [YBB^{+}21]. Render [KKS^{+}20, WSX^{+}22]. Rendering [FHSS18, LY19, MDZV18, PLRD21, RPHD20, TY18, dDB22]. Replate [CSN18]. Resolution [LD18, LB19, LMSS18, RPP21, WSX^{+}22, WXCT19, vdLSE20]. Reuse [KKT^{+}18]. Reversal [Day21]. Revisiting [KKI^{+}18]. Rig [Muk18]. Rigid [LAJB18, BBGB20, Muk18]. Robotic [Far21]. Robotics [BDWG21]. Robust [JFS^{+}21, MEM^{+}20]. Rods [ARM^{+}19]. Rotation [DXS21].

Sample [WS18]. Samples [WS18]. Sampling [EK18, KKS^{+}20, KW21, Muk18, PD19, XO21, XOKN20]. Sampling-based [Muk18]. Scatter [MYS20]. scattering [XOKN20]. Scene [JWB^{+}21, LBR^{+}18, RPP21]. Scenes [vdLSE20]. Scheme [LMSS18]. Screen [TY18, SBE22]. Screen-Space [TY18, SBE22]. Sculpting [MCC22]. Second [NZT19]. Second-Order [NZT19]. Seepage [WFM21]. Selection [LBG18]. Self [TLTM18]. Self-Collision [TLTM18]. Semantic [RPHD20, ZRL21]. Serendipitous [NF21]. Setup [MUEM22]. Shading [MYS20, YZK^{+}19]. Shadows [XJZ^{+}22]. Shan [HCR^{+}21]. Shape [MWW18]. Shapes [LB19]. Sickness [SLW^{+}21, WSX^{+}22]. Signed [MEM^{+}20]. Simplicial [CDGB19]. Simplification [LBR^{+}18]. Simulatable [QLZF21]. Simulation [RHH^{+}22, SLP21, TKC21, WDG^{+}19, WFM21, YCL^{+}19, LRC^{+}22]. Simulator [WSX^{+}22]. single [XOKN20]. skeleton [MMMC21]. Sketch [SDY^{+}18]. Sketch-Based [SDY^{+}18]. Sketching [DAI^{+}18]. Skin [PHM^{+}21, SLP21]. Skinning [KB18]. Soft [CO18, SLP21, XJZ^{+}22]. Soft-Tissue [CO18]. Software [KKT^{+}18]. Soil [WFM21]. Soil-Structure [WFM21]. Solids [KBFF^{+}21]. Solver [AHGB19, NZT19]. Sounds [Rüs21]. Space [TY18, SBE22]. Sparse [HHCM21, STK20]. Spatial [PJL21, TLTM18]. Spatial-Temporal [PJL21]. Speak [Far21]. Speech [FLS^{+}22].

Speech-Driven [FLS^{+}22]. SPH [HRZ^{+}19, KBFF^{+}21]. SPH-centric [HRZ^{+}19]. Sphere [YCL^{+}19]. Spherical [PD19]. Spinning [MD21]. Splatting [FHSS18, SGH18]. Split [LVY^{+}20]. Splitting [EK18]. Stable [XO21]. Stereo [SBE22]. Stereo-consistent [SBE22]. Stereoscopic [CNI^{+}20]. Stochastic [LY20, VSE21]. Stochastic-Depth [VSE21]. Stroke [LZY21]. Structure [RHH^{+}22, WFM21]. Study [SLW^{+}21]. Style [MSK22, PJL21, SCNW19, TTK^{+}21]. Stylization [PJL21]. Sub [GBW20, TKC21]. Sub-Pixel [TKC21]. Sub-triangle [GBW20]. Subaltern [Far21]. Subdivision [LMSS18]. Subsurface [XO21, XOKN20]. Super [WXCT19]. Super-Resolution [WXCT19]. Surface [CWZ^{+}18, YCL^{+}19]. Surfaces [LMSS18, STSK20, WS18]. SYNC [RW21]. Synchronization [WSG19]. Synthesis [CNI^{+}20, PLRD21, STSK20]. Synthesizing [JWB^{+}21]. System [FXX^{+}22, Sim21, WH21].

Taoist [HCR^{+}21]. Technologies [MD21]. Technology [HCR^{+}21]. Temporal [PJL21, SPD18]. Temporally [XO21]. Testing [Yuk18]. Text [FLS^{+}22]. Texture [STSK20, WJG^{+}21]. Thin [WMZ^{+}20]. Three [JDZD19, QLZF21, STSK20]. Three-Level [STSK20]. Three-party [JDZD19]. Tile [BWL18]. Tile-based [BWL18]. Tiled [FHSS18]. Tiles [MYS20]. Tiling [Kir18]. Time [BBGB20, CKY^{+}22, CLS^{+}21, CAS22, LY19, LY20, MUEM22, MSK22, PD19, XJZ^{+}22, YCL^{+}19, CNI^{+}20, KKSM18, LD18, LPW20, SPD18, SCNW19, TTK^{+}21, WH21, XOKN20, XO21, ZCL18]. Time-independent [BBGB20]. Tissue [CO18, SLP21]. Topology [BWL18]. Traced [XJZ^{+}22]. Tracing [HMES20, HHCM21, LMSS18, GBW20]. Tracking [BDWG21]. Tradition [MD21].

Traditional [HCR^{+}21, MD21]. Training [CKY^{+}22]. Trajectory [DMH^{+}21]. Transfer [LWM19, SCNW19, TTK^{+}21, MMMC21]. Transformations [MSK22]. Transformed [KHDN22]. Transforms [WMZ^{+}20]. Translation [SLY20]. Transparency [MKKP18]. transparent [GBW20]. Tree [EK18]. Trees [Dup20, LVY^{+}20, QLZF21]. Trevor [Day21]. triangle [GBW20]. Two [MWW18, SSH^{+}21]. Two-Player [SSH^{+}21]. Two-Way [MWW18]. Types [WMZ^{+}20].

Understanding [KKI^{+}18]. Unified [CDGB19]. UpFlow [RPP21]. Using [JWB^{+}21, WMZ^{+}20, YBB^{+}21, DAI^{+}18, HN18, RPHD20, SGH18, YW19]. Utilizing [TKC21].

Validated [DMH^{+}21]. variance [XOKN20]. variance-guided [XOKN20]. Variates [XO21]. Various [FJdNJ22]. Vertex [KKT^{+}18, KKI^{+}18, PKM22]. Vertex-Blend [PKM22]. via [MSK22, PJL21, STK20, SLY20]. Videos [TTK^{+}21]. View [CNI^{+}20, DAI^{+}18, LBR^{+}18, LBG18, PLRD21]. View-Dependent [LBG18]. View-Region [LBR^{+}18]. VIPER [ARM^{+}19]. Virtual [HMN18, MCC22, SLY20, SLW^{+}21, WSX^{+}22, WH21, CASR21]. Viscous [AHGB19]. Visibility [HMN18, KW21]. Vision [BDWG21, JFS^{+}21, Sim21]. Vision-Based [JFS^{+}21]. Visual [ANEK21, CKY^{+}22, WFM21]. Visualization [LZY21, WDG^{+}19]. Visually [YZK^{+}19]. volcanic [LRC^{+}22]. Volume [ARM^{+}19, SLP21]. Volumes [HHCM21]. Volumetric [DAI^{+}18, HMES20, Wan18]. Voxel [vdLSE20]. VPLs [SGH18]. VR [FXX^{+}22].

Walking [SHS^{+}21]. Warp [KKS^{+}20]. Waterdrop [Fuj21]. Way [MWW18].

Weaving [MD21]. Weights [PHM^{+}21]. Wise [MSK22]. Woods [SSH^{+}21]. Work [Day21].

Zero [KBFF^{+}21]. Zero-Energy [KBFF^{+}21].

References

Aanjaneya:2019:EGM

[AHGB19] Mridul Aanjaneya, Chengguizi Han, Ryan Goldade, and Christopher Batty. An efficient geometric multigrid solver for viscous liquids. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(2):14:1–14:21, July 2019. CODEN ???? ISSN 2577-6193 (electronic). URL

                                                   https://dl.acm.org/doi/10.
                                                   1145/3340255.

Andersson:2020:FDE

[ANAM^{+}20] Pontus Andersson, Jim Nilsson, Tomas Akenine-Möller, Magnus Oskarsson, Kalle Åström, and Mark D. Fairchild. FLIP: a difference evaluator for alternating images. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(2): 15:1–15:23, August 2020. CO-DEN ???? ISSN 2577-6193 (elec-

                                                   tronic). URL https://dl.acm.
                                                   org/doi/10.1145/3406183.

Andrews:2021:CFV

[ANEK21] Sheldon Andrews, Loic Nassif, Kenny Erleben, and Paul G. Kry. Coupling friction with visual appearance. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(3):31:1–31:20, September 2021. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3480138.

Angles:2019:VVI

[ARM^{+}19] Baptiste Angles, Daniel Rebain, Miles Macklin, Brian Wyvill, Loic Barthe, Jp Lewis, Javier Von Der Pahlen, Shahram Izadi, Julien Valentin, Sofien Bouaziz, and Andrea Tagliasacchi. VIPER: Volume invariant position-based elastic rods. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(2): 19:1–19:26, July 2019. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3340260.

Brunel:2020:TID

[BBGB20] Camille Brunel, Pierre Bénard, Gaël Guennebaud, and Pascal Barla. A time-independent deformer for elastic-rigid contacts. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(1):9:1–9:21, April 2020. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3384539.

Bugdayci:2021:IVE

[BDWG21] Irem Bugdayci, Anne-Heloise Dautel, Robert Wuss, and Ruairi Glynn. Instruments of vision: Eye-tracking and robotics as an embodied interface. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(2): 21:1–21:10, July 2021. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3465618.

Bian:2018:TBP

[BWL18] Xiaojun Bian, Li-Yi Wei, and Sylvain Lefebvre. Tile-based pattern design with topology control. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 1(1):23:1–23:15, July 2018. CODEN ???? ISSN 2577-6193 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3203204.

Currius:2022:RTH

[CAS22] Roc R. Currius, Ulf Assarsson, and Erik Sintorn. Realtime hair filtering with convolutional neural networks. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(1): 15:1–15:15, May 2022. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3522606.

Chemistruck:2021:EAP

[CASR21] Mike Chemistruck, Andrew Allen, John Snyder, and Nikunj Raghuvanshi. Efficient acoustic perception for virtual AI agents. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(3):43:1–43:13, September 2021. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3480139.

Chang:2019:USM

[CDGB19] Jumyung Chang, Fang Da, Eitan Grinspun, and Christopher Batty. A unified simplicial model for mixed-dimensional and nonmanifold deformable elastic objects. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(2):11:1–11:18, July 2019. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3340252.

Cardoso:2022:TPV

[CKY^{+}22] Joao Liborio Cardoso, Bernhard Kerbl, Lei Yang, Yury Uralsky, and Michael Wimmer. Training and predicting visual error for real-time applications. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(1): 11:1–11:17, May 2022. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3522625.

Chermain:2021:RTG

[CLS^{+}21] Xavier Chermain, Simon Lucas, Basile Sauvage, Jean-Michel Dischler, and Carsten Dachsbacher. Real-time geometric glint antialiasing with normal map filtering. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(1):1:1–1:16, April 2021. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3451257.

Chaurasia:2020:PRT

[CNI^{+}20] Gaurav Chaurasia, Arthur Nieuwoudt, Alexandru-Eugen Ichim, Richard Szeliski, and Alexander Sorkine-Hornung. Passthrough+: Real-time stereoscopic view synthesis for mobile mixed reality. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(1):7:1– 7:17, April 2020. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3384540.

Casas:2018:LNS

[CO18] Dan Casas and Miguel A. Otaduy. Learning nonlinear soft-tissue dynamics for interactive avatars. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1):10:1–10:15, July 2018. CODEN ???? ISSN 2577-6193 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3203187.

Chen:2018:R

[CSN18] Ge Chen, Pedro V. Sander, and Diego Nehab. The replate. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1): 4:1–4:14, July 2018. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3203205.

Chen:2018:DSL

[CWZ^{+}18] Anpei Chen, Minye Wu, Yingliang Zhang, Nianyi Li, Jie Lu, Shenghua Gao, and Jingyi Yu. Deep surface light fields. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1): 14:1–14:17, July 2018. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3203192.

Delanoy:2018:SUM

[DAI^{+}18] Johanna Delanoy, Mathieu Aubry, Phillip Isola, Alexei A. Efros, and Adrien Bousseau. 3D sketching using multi-view deep volumetric prediction. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1): 21:1–21:22, July 2018. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3203197.

Day:2021:RDP

[Day21] Kevin Day. Reversal, disconnect, and proposition: Noise and data politics in the work of Julian Oliver and Trevor Paglen. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(2): 25:1–25:8, July 2021. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3465624.

deDinechin:2022:RLM

[dDB22] Heloise de Dinechin and Laurent Belcour. Rendering layered materials with diffuse interfaces. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(1):13:1–13:12, May 2022. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3522620.

Daniel:2021:PVM

[DMH^{+}21] Beatríz Cabrero Daniel, Ricardo Marques, Ludovic Hoyet, Julien Pettré, and Josep Blat. A perceptually-validated metric for crowd trajectory quality evaluation. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(3):42:1–42:18, September 2021. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3480136.

Dupuy:2020:CBT

[Dup20] Jonathan Dupuy. Concurrent binary trees (with application to longest edge bisection). Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(2): 21:1–21:20, August 2020. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3406186.

Dong:2021:EAR

[DXS21] Yihong Dong, Lunchen Xie, and Qingjiang Shi. Efficient algorithms for rotation averaging problems. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(1):16:1–16:16, April 2021. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3451263.

Estevez:2018:ISM

[EK18] Alejandro Conty Estevez and Christopher Kulla. Importance sampling of many lights with adaptive tree splitting. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(2): 25:1–25:17, August 2018. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3233305.

Farahi:2021:CSS

[Far21] Behnaz Farahi. Can the subaltern speak?: Feminism in robotic mask design. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(2): 18:1–18:11, July 2021. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3465621.

Franke:2018:MLD

[FHSS18] Linus Franke, Nikolai Hofmann, Marc Stamminger, and Kai Selgrad. Multi-layer depth of field rendering with tiled splatting. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1):6:1–6:17, July 2018. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3203200.

Faria:2022:IPV

[FJdNJ22] Matheus Prado Prandini Faria, Etienne Silva Julia, Marcelo Zanchetta do Nascimento, and Rita Maria Silva Julia. Investigating the performance of various deep neural networks-based approaches designed to identify game events in gameplay footage. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(1):8:1–8:17, May 2022. CODEN ???? ISSN 2577-6193 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3522624.

Fan:2022:JAT

[FLS^{+}22] Yingruo Fan, Zhaojiang Lin, Jun Saito, Wenping Wang, and Taku Komura. Joint audiotext model for expressive speechdriven 3D facial animation. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(1): 16:1–16:15, May 2022. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3522615.

Fujioka:2021:DIA

[Fuj21] Sadam Fujioka. drop: an interactive art installation with waterdrop projection-mapping. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(2):27:1–27:8, July 2021. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3465613.

Fu:2022:EEC

[FXX^{+}22] Zhiying Fu, Rui Xu, Shiqing Xin, Shuangmin Chen, Changhe Tu, Chenglei Yang, and Lin Lu. EasyVRModeling: Easily create 3D models by an immersive VR system. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(1): 10:1–10:14, May 2022. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3522613.

Gruen:2020:STO

[GBW20] Holger Gruen, Carsten Benthin, and Sven Woop. Sub-triangle opacity masks for faster ray tracing of transparent objects. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(2): 18:1–18:12, August 2020. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3406180.

Han:2018:DB

[HCNS18] Songfang Han, Ge Chen, Diego Nehab, and Pedro V. Sander. In-depth buffers. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1):2:1–2:14, July 2018. CODEN ???? ISSN 2577-6193 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3203194.

Henchoz:2021:MSD

[HCR^{+}21] Nicolas Henchoz, Margaux Charvolin, Delphine Ribes, Lara Défayes, Cédric DuchÊne, Emily Groves, and Andreas Sonderegger. Ming Shan digital experience: Immersive technology for traditional Taoist meditation. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(2):24:1–24:10, July 2021. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3465620.

Hildreth:2019:CMA

[HG19] Dalton Hildreth and Stephen J. Guy. Coordinating multi-agent navigation by learning communication. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 2(2):20:1–20:17, July 2019. CODEN ???? ISSN 2577-6193 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3340261.

Han:2019:HMP

[HGG^{+}19] Xuchen Han, Theodore F. Gast, Qi Guo, Stephanie Wang, Chenfanfu Jiang, and Joseph Teran. A hybrid material point method for frictional contact with diverse materials. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(2):17:1–17:24, July 2019. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3340258.

Hofmann:2021:IPT

[HHCM21] Nikolai Hofmann, Jon Hasselgren, Petrik Clarberg, and Jacob Munkberg. Interactive path tracing and reconstruction of sparse volumes. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(1): 5:1–5:19, April 2021. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3451256.

Hofmann:2020:NDP

[HMES20] Nikolai Hofmann, Jana Martschinke, Klaus Engel, and Marc Stamminger. Neural denoising for path tracing of medical volumetric data. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 3(2):13:1–13:18, August 2020. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3406181.

Hunt:2018:HVV

[HMN18] Warren Hunt, Michael Mara, and Alex Nankervis. Hierarchical visibility for virtual reality. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1):8:1–8:18, July 2018. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3203191.

Heitz:2018:HPE

[HN18] Eric Heitz and Fabrice Neyret. High-performance by-example noise using a histogram-preserving blending operator. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(2): 31:1–31:25, August 2018. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3233304.

Huang:2019:GNP

[HRZ^{+}19] Kemeng Huang, Jiming Ruan, Zipeng Zhao, Chen Li, Changbo Wang, and Hong Qin. A general novel parallel framework for SPH-centric algorithms. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(1): 7:1–7:16, June 2019. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3321360.

Jin:2019:DLB

[JDZD19] Aobo Jin, Qixin Deng, Yuting Zhang, and Zhigang Deng. A deep learning-based model for head and eye motion generation in three-party conversations. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(2):9:1–9:19, July 2019. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3340250.

Jonnalagadda:2021:RVB

[JFS^{+}21] Aditya Jonnalagadda, Iuri Frosio, Seth Schneider, Morgan McGuire, and Joohwan Kim. Robust vision-based cheat detection in competitive gaming. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(1): 7:1–7:18, April 2021. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3451259.

Jiang:2021:SIS

[JWB^{+}21] Hao Jiang, Siqi Wang, Huikun Bi, Xiaolei Lv, Binqiang Zhao, Zheng Wang, and Zhaoqi Wang. Synthesizing indoor scene layouts in complicated architecture using dynamic convolution networks. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 4(1):13:1–13:16, April 2021. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3451267.

Komaritzan:2018:PS

[KB18] Martin Komaritzan and Mario Botsch. Projective skinning. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1):12:1–12:19, July 2018. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3203203.

Kugelstadt:2021:FCE

[KBFF^{+}21] Tassilo Kugelstadt, Jan Bender, José Antonio Fernández-Fernández, Stefan Rhys Jeske, Fabian Löschner, and Andreas Longva. Fast corotated elastic SPH solids with implicit zero-energy mode control. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4 (3):33:1–33:21, September 2021. CODEN ???? ISSN 2577-6193 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3480142.

KT:2022:BLT

[KHDN22] Aakash KT, Eric Heitz, Jonathan Dupuy, and P. J. Narayanan. Bringing linearly transformed cosines to anisotropic GGX. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(1):12:1–12:18, May 2022. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3522612.

Kirillov:2018:NPT

[Kir18] Aleksandr Kirillov. Non-periodic tiling of procedural noise functions. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 1(2):32:1–32:15, August 2018. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3233306.

Kadlecek:2019:BAP

[KK19] Petr Kadlecek and Ladislav Kavan. Building accurate physicsbased face models from data. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(2):15:1–15:16, July 2019. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3340256.

Kerbl:2018:RVC

[KKI^{+}18] Bernhard Kerbl, Michael Kenzel, Elena Ivanchenko, Dieter Schmalstieg, and Markus Steinberger. Revisiting the vertex cache: Understanding and optimizing vertex processing on the modern GPU. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(2): 29:1–29:16, August 2018. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3233302.

Kim:2020:PRW

[KKS^{+}20] Joohwan Kim, Pyarelal Knowles, Josef Spjut, Ben Boudaoud, and Morgan Mcguire. Postrender warp with late input sampling improves aiming under high latency conditions. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(2): 12:1–12:18, August 2020. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3406187.

Koniaris:2018:GAD

[KKSM18] Babis Koniaris, Maggie Kosek, David Sinclair, and Kenny Mitchell. GPU-accelerated depth codec for real-time, highquality light field reconstruction. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1):3:1–3:15, July 2018. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3203193.

Kenzel:2018:FVR

[KKT^{+}18] Michael Kenzel, Bernhard Kerbl, Wolfgang Tatzgern, Elena Ivanchenko, Dieter Schmalstieg, and Markus Steinberger. On-the-fly vertex reuse for massively-parallel software geometry processing. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(2): 28:1–28:17, August 2018. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3233303.

Kim:2021:FMO

[KLXvdP21] Nam Hee Kim, Hung Yu Ling, Zhaoming Xie, and Michiel van de Panne. Flexible motion optimization with modulated assistive forces. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(3):35:1–35:25, September 2021. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3480144.

Koch:2021:GVS

[KW21] Thomas Koch and Michael Wimmer. Guided visibility sampling++. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(1):4:1–4:16, April 2021. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3451266.

Li:2018:ERB

[LAJB18] Xiaokai Li, Sheldon Andrews, Ben Jones, and Adam Bargteil. Energized rigid body fracture. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1):9:1–9:9, July 2018. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3203207.

Li:2019:MRM

[LB19] Yijing Li and Jernej Barbic.

Multi-resolution modeling of shapes in contact. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(2):12:1–12:26, July 2019. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3340253.

Lambert:2018:VDM

[LBG18] Thibaud Lambert, Pierre Bénard, and Gaël Guennebaud. A viewdependent metric for patchbased LOD generation 8 selection. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1):20:1–20:21, July 2018. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3203195.

Lall:2018:VRO

[LBR^{+}18] Puneet Lall, Silviu Borac, Dave Richardson, Matt Pharr, and Manfred Ernst. Viewregion optimized image-based scene simplification. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(2): 26:1–26:22, August 2018. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3233311.

Lee:2018:RTA

[LD18] Haebom Lee and Piotr Didyk. Real-time apparent resolution enhancement for head-mounted displays. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1):19:1–19:15, July 2018. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3203202.

Lier:2018:HRC

[LMSS18] Alexander Lier, Magdalena Martinek, Marc Stamminger, and Kai Selgrad. A highresolution compression scheme for ray tracing subdivision surfaces with displacement. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(2): 33:1–33:17, August 2018. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3233308.

Luksch:2020:RTA

[LPW20] Christian Luksch, Lukas Prost, and Michael Wimmer. Realtime approximation of photometric polygonal lights. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(1): 4:1–4:18, April 2020. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3384537.

Lastic:2022:ISP

[LRC^{+}22] Maud Lastic, Damien Rohmer, Guillaume Cordonnier, Claude Jaupart, Fabrice Neyret, and Marie-Paule Cani. Interactive simulation of plume and pyroclastic volcanic ejections. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(1): 4:1–4:15, May 2022. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3522609.

Lin:2020:HAD

[LVY^{+}20] Daqi Lin, Elena Vasiou, Cem Yuksel, Daniel Kopta, and Erik Brunvand. Hardwareaccelerated dual-split trees. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(2): 20:1–20:21, August 2020. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3406185.

Li:2019:DPR

[LWM19] Yue Li, Pablo Wiedemann, and Kenny Mitchell. Deep precomputed radiance transfer for deformable objects. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(1):3:1– 3:16, June 2019. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3320284.

Lin:2019:RTR

[LY19] Daqi Lin and Cem Yuksel. Real-time rendering with lighting grid hierarchy. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(1):8:1– 8:17, June 2019. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3321361.

Lin:2020:RTS

[LY20] Daqi Lin and Cem Yuksel. Realtime stochastic lightcuts. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(1): 5:1–5:18, April 2020. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3384543.

Lyu:2021:ISM

[LZY21] Ruimin Lyu, Tianqin Zhang, and Zhaolin Yuan. Imaginary stroke movement measurement and visualization. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(2): 20:1–20:12, July 2021. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3465625.

Mandal:2022:IPB

[MCC22] Avirup Mandal, Parag Chaudhuri, and Subhasis Chaudhuri. Interactive physics-based virtual sculpting with haptic feedback. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(1):9:1–9:20, May 2022. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3522611.

Montero:2021:SCI

[MD21] Valentina Montero and Sandra De Berduccy. Spinning the conductors of an indigenous tradition: Toward a continuity of traditional Andean weaving with new electronic technologies. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(2): 26:1–26:12, July 2021. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3465622.

Meng:2018:KFR

[MDZV18] Xiaoxu Meng, Ruofei Du, Matthias Zwicker, and Amitabh Varshney. Kernel foveated rendering. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 1(1):5:1–5:20, July 2018. CODEN ???? ISSN 2577-6193 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3203199.

Macklin:2020:LOR

[MEM^{+}20] Miles Macklin, Kenny Erleben, Matthias Müller, Nuttapong Chentanez, Stefan Jeschke, and Zach Corse. Local optimization for robust signed distance field collision. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(1):8:1–8:17, April 2020. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3384538.

Munstermann:2018:MBO

[MKKP18] Cedrick Münstermann, Stefan Krumpen, Reinhard Klein, and Christoph Peters. Momentbased order-independent transparency. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1):7:1–7:20, July 2018. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3203206.

Musoni:2021:FST

[MMMC21] Pietro Musoni, Riccardo Marin, Simone Melzi, and Umberto Castellani. A functional skeleton transfer. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(3):34:1–34:15, September 2021. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3480140.

Mason:2022:RTS

[MSK22] Ian Mason, Sebastian Starke, and Taku Komura. Realtime style modelling of human locomotion via featurewise transformations and local motion phases. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(1): 6:1–6:18, May 2022. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3522618.

Macek:2022:RTR

[MUEM22] Nejc Macek, Baran Usta, Elmar Eisemann, and Ricardo Marroquim. Real-time relighting of human faces with a low-cost setup. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(1):14:1–14:19, May 2022. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3522626.

Mukai:2018:SBR

[Muk18] Tomohiko Mukai. Samplingbased rig conversion into nonrigid helper bones. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1): 13:1–13:17, July 2018. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3203190.

Mukherjee:2018:ITW

[MWW18] Rajaditya Mukherjee, Longhua Wu, and Huamin Wang. Interactive two-way shape design of elastic bodies. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1): 11:1–11:17, July 2018. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3203196.

Mallett:2020:EAD

[MYS20] Ian Mallett, Cem Yuksel, and Larry Seiler. Efficient adaptive deferred shading with hardware scatter tiles. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(2): 11:1–11:17, August 2020. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3406184.

Niehaus:2021:MBA

[NF21] Kiona Hagen Niehaus and Rebecca Fiebrink. Making up 3D bodies: Artistic and serendipitous modeling of digital human figures. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(2):23:1–23:9, July 2021. CODEN ???? ISSN 2577-6193 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3468779.

Narain:2019:SOA

[NZT19] Rahul Narain, Jonas Zehnder, and Bernhard Thomaszewski. A second-order advection-reflection solver. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 2(2):16:1–16:14, July 2019. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3340257.

Ogaki:2020:GLP

[Oga20] Shinji Ogaki. Generalized light portals. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 3(2):10:1–10:19, August 2020. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3406176.

Peters:2019:SPS

[PD19] Christoph Peters and Carsten Dachsbacher. Sampling projected spherical caps in real time. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(1):1:1–1:16, June 2019. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3320282.

Pan:2021:HHN

[PHM^{+}21] Xiaoyu Pan, Jiancong Huang, Jiaming Mai, He Wang, Honglin Li, Tongkui Su, Wenjun Wang, and Xiaogang Jin. HeterSkinNet: a heterogeneous network for skin weights prediction. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(1): 10:1–10:19, April 2021. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3451262.

Park:2021:DMS

[PJL21] Soomin Park, Deok-Kyeong Jang, and Sung-Hee Lee. Diverse motion stylization for multiple style domains via spatialtemporal graph-based generative model. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(3):36:1–36:17, September 2021. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3480145.

Peters:2022:PCV

[PKM22] Christoph Peters, Bastian Kuth, and Quirin Meyer. Permutation coding for vertex-blend attribute compression. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(1): 5:1–5:16, May 2022. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3522607.

Prakash:2021:HIB

[PLRD21] Siddhant Prakash, Thomas Leimkühler, Simon Rodriguez, and George Drettakis. Hybrid image-based rendering for freeview synthesis. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(1):15:1–15:20, April 2021. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3451260.

Quigley:2021:TDR

[QLZF21] Ed Quigley, Winnie Lin, Yilin Zhu, and Ronald Fedkiw. Three dimensional reconstruction of botanical trees with simulatable geometry. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(3):37:1–37:16, September 2021. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3480146.

Raateland:2022:DAG

[RHH^{+}22] Wouter Raateland, Torsten Hädrich, Jorge Alejandro Amador Herrera, Daniel T. Banuti, Wojciech Pałubicki, Sören Pirk, Klaus Hildebrandt, and Dominik L. Michels. DCGrid: an adaptive grid structure for memory-constrained fluid simulation on the GPU. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(1): 3:1–3:14, May 2022. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3522608.

Reshetov:2018:PRH

[RL18] Alexander Reshetov and David Luebke. Phantom ray-hair intersector. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 1(2):34:1–34:22, August 2018. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3233307.

Rodriguez:2020:IBR

[RPHD20] Simon Rodriguez, Siddhant Prakash, Peter Hedman, and George Drettakis. Image-based rendering of cars using semantic labels and approximate reflection flow. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(1):6:1–6:17, April 2020. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3384535.

Roy:2021:NUS

[RPP21] Bruno Roy, Pierre Poulin, and Eric Paquette. Neural Up-Flow: a scene flow learning approach to increase the apparent resolution of particlebased liquids. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(3):40:1–40:26, September 2021. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3480147.

Rust:2021:BMG

[Rüs21] Annina Rüst. Bad mother /good mother: The poetics and politics of the sounds of invisible labor. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(2):17:1–17:9, July 2021. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3465617.

Riahi:2021:PBN

[RW21] Maryam Riahi and Benjamin Allen Watson. Am I playing better now?: The effects of G-SYNC in 60Hz gameplay. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(1): 6:1–6:17, April 2021. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3451269.

Shi:2022:SCS

[SBE22] Peiteng Shi, Markus Billeter, and Elmar Eisemann. Stereoconsistent screen-space ambient occlusion. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(1):2:1–2:12, May 2022. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3522614.

Smith:2019:ENN

[SCNW19] Harrison Jesse Smith, Chen Cao, Michael Neff, and Yingying Wang. Efficient neural networks for real-time motion style transfer. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(2):13:1–13:17, July 2019. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3340254.

Su:2018:ISB

[SDY^{+}18] Wanchao Su, Dong Du, Xin Yang, Shizhe Zhou, and Hongbo Fu. Interactive sketch-based normal map generation with deep neural networks. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1): 22:1–22:17, July 2018. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3203186.

Sriwasansak:2018:EEC

[SGH18] Jamorn Sriwasansak, Adrien Gruson, and Toshiya Hachisuka. Efficient energy-compensated VPLs using photon splatting. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1):16:1–16:13, July 2018. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3203189.

Schmelter:2021:IBR

[SHS^{+}21] Thereza Schmelter, Levente Hernadi, Marc Aurel Störmer, Frank Steinicke, and Kristian Hildebrand. Interaction based redirected walking. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(1):9:1–9:16, April 2021. CODEN ???? ISSN 2577-6193 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3451264.

Simon:2021:JVE

[Sim21] Gilles Simon. Jan van Eyck’s perspectival system elucidated through computer vision. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(2): 22:1–22:8, July 2021. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3465623.

Sheen:2021:VPS

[SLP21] Seung Heon Sheen, Egor Larionov, and Dinesh K. Pai. Volume preserving simulation of soft tissue with skin. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4 (3):32:1–32:23, September 2021. CODEN ???? ISSN 2577-6193 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3480143.

Shi:2021:VRS

[SLW^{+}21] Rongkai Shi, Hai-Ning Liang, Yu Wu, Difeng Yu, and Wenge Xu. Virtual reality sickness mitigation methods: a comparative study in a racing game. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(1): 8:1–8:16, April 2021. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3451255.

Seiler:2020:CCG

[SLY20] Larry Seiler, Daqi Lin, and Cem Yuksel. Compacted CPU/GPU data compression via modified virtual address translation. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(2): 19:1–19:18, August 2020. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3406177.

Schied:2018:GER

[SPD18] Christoph Schied, Christoph Peters, and Carsten Dachsbacher. Gradient estimation for realtime adaptive temporal filtering. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 1(2):24:1–24:16, August 2018. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3233301.

Schreiner:2021:GPP

[SPL^{+}21] Paul Schreiner, Maksym Perepichka, Hayden Lewis, Sune Darkner, Paul G. Kry, Kenny Erleben, and Victor B. Zordan. Global position prediction for interactive motion capture. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4 (3):39:1–39:16, September 2021. CODEN ???? ISSN 2577-6193 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3479985.

Swearingen:2021:WMR

[SSH^{+}21] Kyoung Swearingen, Scott Swearingen, Fede Camara Halac, Sruthi Ammannagari, and Matt Hall. The Woods: a mixed-reality two-player cooperative game. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(2):30:1–30:7, July 2021. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3465616.

Schuster:2020:HPI

[STK20] Kersten Schuster, Philip Trettner, and Leif Kobbelt. Highperformance image filters via sparse approximations. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(2): 14:1–14:19, August 2020. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3406182.

Schuster:2020:TLA

[STSK20] Kersten Schuster, Philip Trettner, Patric Schmitz, and Leif Kobbelt. A three-level approach to texture mapping and synthesis on 3D surfaces. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(1): 1:1–1:19, April 2020. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3384542.

Tumanov:2021:DDP

[TKC21] Evgenii Tumanov, Dmitry Korobchenko, and Nuttapong Chentanez. Data-driven particlebased liquid simulation with deep learning utilizing sub-pixel convolution. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(1):12:1–12:16, April 2021. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3451261.

Tang:2018:PPS

[TLTM18] Min Tang, Zhongyuan Liu, Ruofeng Tong, and Dinesh Manocha. PSCC: Parallel selfcollision culling with spatial hashing on GPUs. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1):18:1–18:18, July 2018. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3203188.

Texler:2021:FIR

[TTK^{+}21] Aneta Texler, Ondrej Texler, Michal Kucera, Menglei Chai, and Daniel Sýkora. FaceBlit: Instant real-time example-based style transfer to facial videos. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(1):14:1–14:17, April 2021. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3451270.

Truong:2018:NRF

[TY18] Nghia Truong and Cem Yuksel. A narrow-range filter for screenspace fluid rendering. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1): 17:1–17:15, July 2018. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3203201.

Truong:2020:QAC

[TYS20] Nghia Truong, Cem Yuksel, and Larry Seiler. Quadratic approximation of cubic curves. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(2): 16:1–16:17, August 2020. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3406178.

vanderLaan:2020:LGC

[vdLSE20] Remi van der Laan, Leonardo Scandolo, and Elmar Eisemann. Lossy geometry compression for high resolution voxel scenes. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(1):2:1–2:13, April 2020. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3384541.

Viitanen:2018:PFH

[VKJ^{+}18] Timo Viitanen, Matias Koskela, Pekka Jääskeläinen, Aleksi Tervo, and Jarmo Takala. PLOCTree: a fast, high-quality hardware BVH builder. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(2): 35:1–35:19, August 2018. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3233309.

Vermeer:2021:SDA

[VSE21] Jop Vermeer, Leonardo Scandolo, and Elmar Eisemann. Stochastic-depth ambient occlusion. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(1):3:1–3:15, April 2021. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3451268.

Wang:2018:FMC

[Wan18] Cuilan Wang. A fast method to calculate the volumetric divergence metric for evaluating the accuracy of the extracted isosurface. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(2): 27:1–27:19, August 2018. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3233312.

Wang:2019:SVD

[WDG^{+}19] Stephanie Wang, Mengyuan Ding, Theodore F. Gast, Leyi Zhu, Steven Gagniere, Chenfanfu Jiang, and Joseph M. Teran. Simulation and visualization of ductile fracture with the material point method. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(2): 18:1–18:20, July 2019. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3340259.

Wang:2021:VSS

[WFM21] Xu Wang, Makoto Fujisawa, and Masahiko Mikawa. Visual simulation of soil-structure destruction with seepage flows. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4 (3):41:1–41:18, September 2021. CODEN ???? ISSN 2577-6193 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3480141.

Wu:2021:MGA

[WH21] Ziwei Wu and Lingdong Huang.

Mimicry: Genetic-algorithm-based real-time system of virtual insects in a living environment — a new and altered nature. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(2):28:1–28:8, July 2021. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3465615.

Wu:2021:RGI

[WJG^{+}21] Jane Wu, Yongxu Jin, Zhenglin Geng, Hui Zhou, and Ronald Fedkiw. Recovering geometric information with learned texture perturbations. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(3):38:1–38:18, September 2021. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3480137.

Wald:2020:UHR

[WMZ^{+}20] Ingo Wald, Nate Morrical, Stefan Zellmann, Lei Ma, Will Usher, Tiejun Huang, and Valerio Pascucci. Using hardware ray transforms to accelerate ray/ primitive intersections for long, thin primitive types. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(2): 17:1–17:16, August 2020. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3406179.

Wang:2018:FGP

[WS18] Tong Wang and Reiji Suda. Fast generation of Poisson-disk samples on mesh surfaces by progressive sample projection. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(2): 30:1–30:18, August 2018. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3233310.

Watson:2019:EAS

[WSG19] Benjamin Watson, Rachit Shrivastava, and Ajinkya Gavane. The effects of adaptive synchronization on performance and experience in gameplay. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(1): 5:1–5:13, June 2019. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3320286.

Wang:2022:ERR

[WSX^{+}22] Jialin Wang, Rongkai Shi, Zehui Xiao, Xueying Qin, and Hai-Ning Liang. Effect of render resolution on gameplay experience, performance, and simulator sickness in virtual reality games. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(1):7:1–7:15, May 2022. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3522610.

Werhahn:2019:MPG

[WXCT19] Maximilian Werhahn, You Xie, Mengyu Chu, and Nils Thuerey. A multi-pass GAN for fluid flow super-resolution. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(2): 10:1–10:21, July 2019. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3340251.

Xu:2022:RTR

[XJZ^{+}22] Yang Xu, Yuanfa Jiang, Junbo Zhang, Kang Li, and Guohua Geng. Real-time ray-traced soft shadows of environmental lighting by conical ray culling. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 5(1): 1:1–1:15, May 2022. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3522617.

Xu:2021:GLA

[XK21] Pei Xu and Ioannis Karamouzas. A GAN-like approach for physics-based imitation learning and interactive character control. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(3):44:1–44:22, September 2021. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3480148.

Xie:2021:RTS

[XO21] Tiantian Xie and Marc Olano. Real-time subsurface control variates: Temporally stable adaptive sampling. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(1):2:1– 2:18, April 2021. CODEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3451265.

Xie:2020:RTS

[XOKN20] Tiantian Xie, Marc Olano, Brian Karis, and Krzysztof Narkowicz. Real-time subsurface scattering with single pass variance-guided adaptive importance sampling. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 3(1):3:1–3:21, April 2020. CO-DEN ???? ISSN 2577-6193 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3384536.

Yi:2021:RUB

[YBB^{+}21] Changyoon Yi, Juhyun Bae, Nakkyu Baek, Jina Jung, Sunwoong Hur, Hyun Jean Lee, and Seung Ah Lee. ReMember: Using biosignals to recall memories of companion animals. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(2): 29:1–29:7, July 2021. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3465614.

Yang:2019:RTF

[YCL^{+}19] Bowen Yang, William Corse, Jiecong Lu, Joshuah Wolper, and Chen-Fanfu Jiang. Realtime fluid simulation on the surface of a sphere. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(1):4:1– 4:17, June 2019. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3320285.

Yuksel:2018:ADA

[Yuk18] Cem Yuksel. Alpha distribution for alpha testing. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 1(1): 1:1–1:11, July 2018. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3203185.

Yang:2019:ABC

[YW19] Hsiang-Yu Yang and Sai-Keung Wong. Agent-based cooperative animation for box-manipulation using reinforcement learning. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(1):2:1–2:18, June 2019. CO-DEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3320283.

Yang:2021:EHO

[YY21] Zeshi Yang and Zhiqi Yin. Efficient hyperparameter optimization for physics-based character animation. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(1):11:1–11:19, April 2021. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3451254.

Yang:2019:VLC

[YZK^{+}19] Lei Yang, Dmitry Zhdan, Emmett Kilgariff, Eric B. Lum, Yubo Zhang, Matthew Johnson, and Henrik Rydgård. Visually lossless content and motion adaptive shading in games. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 2(1): 6:1–6:19, June 2019. CODEN ???? ISSN 2577-6193 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3320287.

Zanni:2018:HHR

[ZCL18] Cédric Zanni, Frédéric Claux, and Sylvain Lefebvre. HCSG: Hashing for real-time CSG modeling. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACM-CGIT), 1(1):15:1–15:19, July 2018. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3203198.

Zhang:2021:CPI

[ZRL21] Weidi Zhang, Donghao Ren, and George Legrady. Cangjie’s poetry: an interactive art experience of a semantic humanmachine reality. Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT), 4(2):19:1–19:9, July 2021. CODEN ???? ISSN 2577-6193 (electronic). URL

https://dl.acm.org/doi/10.
1145/3465619.